# A Starting Point for Dynamic Community Detection with Leiden Algorithm

### v1

@@ Abstract.
Many real-world graphs evolve with time. Identifying communities or clusters on such graphs is an important problem. In this technical report, we extend three dynamic approaches, namely, Naive-dynamic (ND), Delta-screening (DS), and Dynamic Frontier (DF), to our multicore implementation of the Leiden algorithm, an algorithm known for its high-quality community detection. Our experiments on a server with a 64-core AMD EPYC-7742 processor demonstrate that ND, DS, and DF Leiden achieve speedups of 1.25×1.25$\times$1.25 ×, 1.24×1.24$\times$1.24 ×, and 1.37×1.37$\times$1.37 × on large graphs with random batch updates, compared to Static, ND, and DS Leiden, respectively. However, on real-world dynamic graphs, ND Leiden performs the best, being on average 1.14×1.14$\times$1.14 × faster than Static Leiden. We hope our early results serve as a starting point for dynamic approaches to the Leiden algorithm on evolving graphs.
@@ 1. Introduction
Identifying hidden communities within networks is a crucial graph analytics problem relevant to various domains, including topic discovery, disease prediction, drug discovery, criminal identification, protein annotation, and inferring land use. The goal is to identify groups of vertices with dense internal connections and sparse connections to the rest of the graph. A challenge in community detection is the absence of prior knowledge about the number and size distribution of communities. To address this, researchers have developed numerous heuristics for finding communities (Guimera and Amaral, 2005; Derényi et al., 2005; Newman, 2006; Reichardt and Bornholdt, 2006; Raghavan et al., 2007; Blondel et al., 2008; Rosvall and Bergstrom, 2008; Rosvall et al., 2009; Fortunato, 2010; Gregory, 2010; Kloster and Gleich, 2014; Côme and Latouche, 2015; Ruan et al., 2015; Newman and Reinert, 2016; Ghoshal et al., 2019; Rita, 2020; Lu and Chakraborty, 2020; Gupta et al., 2022). The quality of identified communities is often measured using fitness metrics such as the modularity score proposed by Newman et al. (Newman, 2006). These communities are intrinsic when identified based solely on network topology, and they are disjoint when each vertex belongs to only one community (Gregory, 2010).
@@
The Louvain method, proposed by Blondel et al. (Blondel et al., 2008) from the University of Louvain, is one of the most popular community detection algorithms (Lancichinetti and Fortunato, 2009). This greedy algorithm employs a two-phase approach, consisting of a local-moving phase and an aggregation phase, to iteratively optimize the modularity metric (Blondel et al., 2008). It has a time complexity of $O(KM)$ (where $M$ is the number of edges in the graph and $K$ is the total number of iterations performed across all passes) and efficiently identifies communities with high modularity.
@@
Despite its popularity, the Louvain method has been observed to produce internally disconnected and poorly connected communities. To address these shortcomings, Traag et al. (Traag et al., 2019), from the University of Leiden, proposed the Leiden algorithm. This algorithm introduces an additional refinement phase between the local-moving and aggregation phases. During the refinement phase, vertices can explore and potentially form sub-communities within the communities identified in the local-moving phase. This allows the Leiden algorithm to identify well-connected communities (Traag et al., 2019).
@@
But many real-world graphs evolve rapidly over time, through the insertion and deletion of edges and vertices. These dynamic graphs are often immense in scale, arising from applications such as machine learning and social networks, and are becoming increasingly ubiquitous. For efficiency, algorithms are needed that update results without recomputing from scratch, known as dynamic algorithms. One straightforward strategy for dynamic community detection involves leveraging the community memberships of vertices from the previous snapshot of the graph (Aynaud and Guillaume, 2010; Chong and Teow, 2013; Shang et al., 2014; Zhuang et al., 2019), which we term as Naive-dynamic (ND). Alternatively, more sophisticated techniques have been devised to reduce computational overhead by identifying a smaller subset of the graph affected by changes. These techniques include updating only changed vertices (Aktunc et al., 2015; Yin et al., 2016), processing vertices in the proximity of updated edges (within a specified threshold distance) (Held et al., 2016), disbanding affected communities into lower-level networks (Cordeiro et al., 2016), or employing a dynamic modularity metric to recompute community memberships from scratch (Meng et al., 2016). Recently, a technique known as Delta-screening (DS) has been proposed, which identifies a subset of vertices impacted by changes in a graph using delta-modularity (Zarayeneh and Kalyanaraman, 2021). We have also introduced an efficient incrementally expanding approach for dynamic community detection, which we call the Dynamic Frontier (DF) approach (Sahu, 2024).
However, the above efforts have focused on detecting communities in dynamic networks using the Louvain algorithm. None have extended these approaches to the Leiden algorithm. Further, a majority of the reported algorithms (Aynaud and Guillaume, 2010; Chong and Teow, 2013; Meng et al., 2016; Cordeiro et al., 2016; Zhuang et al., 2019; Zarayeneh and Kalyanaraman, 2021) are sequential. Parallel algorithms for graph analytics on dynamic graphs are an active area of research. Examples of parallel dynamic algorithms include those for updating centrality scores (Shao et al., 2020; Regunta et al., 2020), maintaining shortest paths (Zhang et al., 2017; Khanda et al., 2021), and dynamic graph coloring (Yuan et al., 2017; Bhattacharya et al., 2018).
This technical report extends the parallel Dynamic Frontier (DF) approach to the Leiden algorithm. [^1] https://github.com/puzzlef/leiden-communities-openmp-dynamic DF Leiden is based on GVE-Leiden (Sahu, 2023), our multicore implementation of the Static Leiden algorithm. Upon receiving a batch update comprising edge deletions and insertions, DF Leiden incrementally identifies and processes an approximate set of affected vertices in an incremental manner. Additionally, we introduce parallel implementations of Naive-dynamic (ND) and Delta-screening (DS) Leiden. This work represents, to the best of our knowledge, the first endeavor in extending existing dynamic approaches to the Leiden algorithm.
@@ 2.1. Community detection
Disjoint community detection involves establishing a community membership mapping, $C:V\rightarrow\Gamma$, which maps each vertex $i\in V$ to a community ID $c\in\Gamma$, where $\Gamma$ is the set of community-ids. The vertices within a community $c\in\Gamma$ are denoted as $V_{c}$, and the community to which a vertex $i$ belongs is denoted as $C_{i}$. Furthermore, we define the neighbors of vertex $i$ belonging to a community $c$ as $J_{i\rightarrow c}=\{j\ |\ j\in J_{i}\ and\ C_{j}=c\}$, the sum of those edge weights as $K_{i\rightarrow c}=\{w_{ij}\ |\ j\in J_{i\rightarrow c}\}$, the sum of weights of edges within a community $c$ as $\sigma_{c}=\sum_{(i,j)\in E\ and\ C_{i}=C_{j}=c}w_{ij}$, and the total edge weight of a community $c$ as $\Sigma_{c}=\sum_{(i,j)\in E\ \mbox{and}\ C_{i}=c}w_{ij}$ (Zarayeneh and Kalyanaraman, 2021; Leskovec, 2021).
@@ 2.2. Modularity
Modularity is a fitness metric used to assess the quality of communities produced by community detection algorithms, which are typically heuristic-based. It is calculated as the difference between the fraction of edges within communities and the expected fraction of edges if they were distributed randomly. Modularity ranges from $[-0.5,1]$ (higher values indicate better quality) (Brandes et al., 2007). Theoretically, optimizing this function leads to the best possible grouping (Newman, 2004; Traag et al., 2011).
The modularity $Q$ of the obtained communities can be calculated using Equation 1, where $\delta$ represents the Kronecker delta function ($\delta(x,y)=1$ if $x=y$ and 00 otherwise). The delta modularity of moving a vertex $i$ from community $d$ to community $c$, denoted as $\Delta Q_{i:d\rightarrow c}$, is determined using Equation 2.
(1) $Q=\frac{1}{2m}\sum_{(i,j)\in E}\left[w_{ij}-\frac{K_{i}K_{j}}{2m}\right]\delta(C_{i},C_{j})=\sum_{c\in\Gamma}\left[\frac{\sigma_{c}}{2m}-\left(\frac{\Sigma_{c}}{2m}\right)^{2}\right]$
@@ 2.3. Louvain algorithm
The Louvain method is a greedy, modularity optimization-based agglomerative algorithm that identifies high-quality communities within a graph (Lancichinetti and Fortunato, 2009). The method operates in two phases. In the local-moving phase, each vertex $i$ evaluates moving to the community of one of its neighbors $j\in J_{i}$ that maximizes the increase in modularity $\Delta Q_{i:C_{i}\rightarrow C_{j}}$ (using Equation 2). In the aggregation phase, all vertices within the same community are collapsed into a super-vertex. These two phases constitute one pass, which is repeated until there is no further gain in modularity. The method produces a hierarchy of community memberships for each vertex, forming a dendrogram, with the top-level representing the final communities (Leskovec, 2021).
@@ 2.5. Dynamic approaches
(a) Naive-dynamic (ND) approach
(b) Delta-screening (DS) approach
(c) Dynamic Frontier (DF) approach
Figure 1. Comparison of dynamic community detection approaches: Naive-dynamic (ND), Delta-screening (DS), and Dynamic Frontier (DF) approaches. Edge deletions and insertions are indicated by dotted lines. Vertices (initially) marked as affected by each approach are highlighted in yellow, and when entire communities are marked as affected, they are hatched (Sahu, 2024).
2.5.1. Naive-dynamic (ND) approach
The Naive-dynamic approach, presented by Aynaud et al. (Aynaud and Guillaume, 2010), is a straightforward method for detecting communities in dynamic networks. In this approach, vertices are assigned to communities based on the previous snapshot of the graph, and all vertices are processed, regardless of the edge deletions and insertions in the batch update (hence the term naive). This is illustrated in Figure 1(a), where all vertices are marked as affected and highlighted in yellow. As all communities are also marked as affected, they are depicted with hatching. In the figure, edge deletions are shown in the top row (indicated by dotted lines), edge insertions are shown in the middle row (also indicated by dotted lines), and vertex migration during the community detection algorithm is shown in the bottom row. The community membership obtained through this method is guaranteed to be at least as accurate as that obtained by the static algorithm.
2.5.2. Delta-screening (DS) approach
@@
The Delta-screening (DS) approach, proposed by Zarayeneh et al. (Zarayeneh and Kalyanaraman, 2021), is a dynamic community detection approach that utilizes modularity-based scoring to identify an approximate region of the graph where vertices are likely to alter their community membership. Figure 1(b) illustrates the vertices (and communities) connected to a single source vertex $i$ that are identified as affected by the DS approach in response to a batch update involving both edge deletions and insertions. In the DS approach, Zarayeneh et al. first sort the batch update consisting of edge deletions $(i,j)\in\Delta^{t-}$ and insertions $(i,j,w)\in\Delta^{t+}$ by their source vertex ID. For edge deletions within the same community, they mark $i$’s neighbors and $j$’s community as affected. For edge insertions across communities, they select the vertex $j^{*}$ with the highest modularity change among all insertions linked to vertex $i$ and mark $i$’s neighbors and $j^{*}$’s community as affected. Edge deletions between different communities and edge insertions within the same community are unlikely to impact community membership and are therefore ignored. We process the vertices marked as affected by the DS approach in the first pass of Leiden, with the community membership of each vertex initialized to that obtained in the previous snapshot of the graph. In subsequent passes, all super-vertices are designated as affected and processed.
2.5.3. Dynamic Frontier (DF) approach
@@
When processing a batch update containing edge deletions $(i,j,w)\in\Delta^{t-}$ and insertions $(i,j,w)\in\Delta^{t+}$ upon the original graph, it is likely that only a small subset of vertices in the graph would undergo a change in their community membership. However, the ND approach processes all vertices, while the DS approach tends to overestimate the set of affected vertices, while incurring a high overhead in estimating the set of affected vertices. To tackle this issue, we recently introduced the Dynamic Frontier (DF) approach (Sahu, 2024). It commences by initializing each vertex’s community membership to that obtained in the previous graph snapshot. Figure 1(c) illustrates the vertices connected to a single source vertex $i$ identified as affected by the DF approach in response to a batch update. In instances where edge deletions occur between vertices within the same community or edge insertions between vertices in different communities, the DF approach marks the source vertex $i$ as affected, as depicted by the yellow-highlighted vertices in Figure 1(c). It is noteworthy that batch updates are undirected, hence both endpoints $i$ and $j$ are effectively marked. Edge deletions spanning different communities or edge insertions within the same community are disregarded (as previously mentioned in Section 2.5.2). Furthermore, when a vertex $i$ alters its community membership during the community detection process (illustrated by $i$ transitioning from its original community in the center to its new community on the right), all its neighboring vertices $j\in J_{i}$ are marked as affected, as depicted in Figure 1(c) (highlighted in yellow), while $i$ is marked as unaffected. To minimize unnecessary computation, an affected vertex $i$ is also marked as unaffected even if its community remains unchanged. This optimization, referred to as vertex pruning (Ozaki et al., 2016; Ryu and Kim, 2016; Shi et al., 2021; Zhang et al., 2021). Thus, the DF approach follows a graph traversal-like process until the vertices’ community assignments converge. Similar to the DS approach, it is applied to the first pass of the Leiden algorithm.

### v4

@@ Abstract.
Real-world graphs often evolve over time, making community or cluster detection a crucial task. In this technical report, we extend three dynamic approaches — Naive-dynamic (ND), Delta-screening (DS), and Dynamic Frontier (DF) — to our multicore implementation of the Leiden algorithm, known for its high-quality community detection. Our experiments, conducted on a server with a 64-core AMD EPYC-7742 processor, show that ND, DS, and DF Leiden achieve average speedups of 1.37×1.37$\times$1.37 ×, 1.47×1.47$\times$1.47 ×, and 1.98×1.98$\times$1.98 × on large graphs with random batch updates, compared to the Static Leiden algorithm — while scaling at a rate of 1.6×1.6$\times$1.6 × for every doubling of threads. To our knowledge, this is the first attempt to apply dynamic approaches to the Leiden algorithm. We hope these early results pave the way for further development of dynamic approaches for evolving graphs.
@@ 1. Introduction
Driven by the ability of graphs to represent complex real-world data and capture intricate relationships among entities, research in graph-structured data has been rapidly growing. A core focus of this field is community detection, which involves decomposing a graph into densely connected groups — revealing the natural structure inherent in the data. Discovering hidden communities in social networks (Blekanov et al., 2021), analyzing regional retail landscapes (Verhetsel et al., 2022), studying user relations in Decentralized Online Social Networks (DOSNs) (La Cava et al., 2022), delineating health care service areas (Wang et al., 2021), uncovering disinformation networks in Telegram (La Morgia et al., 2021), partitioning large graphs for machine learning (Bai et al., 2024), studying biological processes (Heumos et al., 2023), characterizing polarized information ecosystems (Uyheng et al., 2021), developing cyber resilient systems (Chernikova et al., 2022), identifying transportation patterns (Chen et al., 2023), identifying attacks in blockchain networks (Erfan et al., 2023), examining restored accounts on Twitter (Kapoor et al., 2021), exploring the eco-epidemiology of zoonoses (Desvars-Larrive et al., 2024), analyzing linguistic variation in memes (Zhou et al., 2023), reconstructing multi-step cyber attacks (Zang et al., 2023), studying twitter communities during the 2022 war in Ukraine (Sliwa et al., 2024), and automated microservice decomposition (Cao and Zhang, 2022) are all applications of community detection.
@@
A challenge in community detection is the absence of prior knowledge about the number and size distribution of communities. To address this, researchers have developed numerous heuristics for finding communities (Blondel et al., 2008; Gregory, 2010). The quality of identified communities is often measured using fitness metrics such as the modularity score proposed by Newman et al. (Newman, 2004). These communities are intrinsic when identified based solely on network topology, and they are disjoint when each vertex belongs to only one community (Gregory, 2010).
@@
The Louvain method, proposed by Blondel et al. (Blondel et al., 2008), is one of the most popular community detection algorithms (Lancichinetti and Fortunato, 2009). This greedy algorithm employs a two-phase approach, consisting of an iterative local-moving phase and an aggregation phase, to iteratively optimize the modularity metric over several passes (Blondel et al., 2008). It has a time complexity of $O(LM)$ (where $M$ is the number of edges in the graph and $L$ is the total number of iterations performed across all passes) and efficiently identifies communities with high modularity.
@@
Despite its popularity, the Louvain method has been observed to produce internally disconnected and poorly connected communities. This problem occurs during the local-moving phase, where important nodes — that serve as bridges within a community — may be reassigned to another community to which they have stronger connections, breaking the connectivity of their original community. To address these issues, Traag et al. (Traag et al., 2019) proposed the Leiden algorithm, which introduces a refinement phase between the local-moving and aggregation phases. During the refinement phase, vertices can explore and potentially form sub-communities within the communities identified in the local-moving phase — allowing the algorithm to identify well-connected communities (Traag et al., 2019).
But many real-world graphs are immense and evolve rapidly over time. For efficiency, algorithms are needed that update results without recomputing from scratch, known as dynamic algorithms. Dynamic community detection algorithms also allow one to track the evolution of communities over time, identifying key events like growth, shrinkage, merging, splitting, birth, and death. However, research efforts have focused on detecting communities in dynamic networks using the Louvain algorithm. None of the works have extended these approaches to the Leiden algorithm.
This technical report extends the Naive-dynamic (ND) (Aynaud and Guillaume, 2010), Delta-screening (DS) (Zarayeneh and Kalyanaraman, 2021), and the recently proposed parallel Dynamic Frontier (DF) approach (Sahu et al., 2024b) to the Leiden algorithm. [^1] https://github.com/puzzlef/leiden-communities-openmp-dynamic Our algorithms build on top of GVE-Leiden (Sahu et al., 2024a), one of the most efficient shared-memory implementations of the Static Leiden algorithm.
1.1. Our Contributions
We propose the first dynamic algorithms for Leiden, extending dynamic approaches originally developed for Louvain (Aynaud and Guillaume, 2010; Zarayeneh and Kalyanaraman, 2021; Sahu et al., 2024b) to Leiden. Our contributions are as follows:
  • Directly applying DF approach (Sahu et al., 2024b) to Leiden fails: On a small batch update Leiden’s local-moving phase converges quickly. Then refinement phase, divides communities into smaller sub-communities. Terminating at this point (as DF Louvain) results in poor modularity, as these sub-communities require hierarchical aggregation. Consequently, multiple passes of Leiden are needed despite early convergence. This is a key insight.
  • For selective refinement we track vertices that migrate between communities during local-moving phase, and marking both source and target communities for refinement. Untouched communities are not refined. This reduces the processing costs, as refined communities must be processed in the subsequent passes.
  • Refinement begins by initializing each vertex as its own sub-community. These merge into larger sub-communities within their own community bounds. However, this approach applied to a subset of communities can result in internal disconnection. This occurs because a vertex $i$ may not belong to community $C$ with the same ID (i.e., $C=i$), but the initialization places $i$ in an isolated sub-community, while other vertices in community $C$ may not be connected to $i$. As refinement continues, more vertices join $i$ in sub-community $D=i$, yet it remains disconnected from rest of community $C=i$. This insight is illustrated in Figure 3, and explained in detail in Section 4.3.
  • To resolve this, we propose a subset renumbering procedure where, for each community $C$, the ID of any member vertex $i^{\prime}$ is selected. This ID is then used to renumber the community membership of all vertices in $C$, update the total edge weight, and set the changed communities flags to the new ID $j$. Afterward, selective refinement can be performed.
  • We also select communities for refinement based on the batch update (changed communities) - those with edge deletions or insertions within the community. Cross-community edge deletions and insertions are automatically selected when vertices migrate between communities.
  • Selective refinement causes load imbalance during aggregation phase, as unrefined communities remain large while refined ones are small. To address this, we reduce chunk size for OpenMP parallelization during aggregation, which introduces scheduling overhead but ensures load balancing.
  • We present novel algorithms, including methods for marking changed-communities and subset-renumbering.
While reported speedups are modest (DF Leiden is 3.7×3.7$\times$3.7 × faster than Static Leiden for small updates), we believe these insights are crucial for developing more efficient dynamic algorithms for Leiden.
2. Related work
@@
A core idea for dynamic community detection, among most approaches, is to use the community membership of each vertex from the previous snapshot of the graph, instead of initializing each vertex into singleton communities. One straightforward strategy for dynamic community detection involves leveraging the community memberships of vertices from the previous snapshot of the graph (Aynaud and Guillaume, 2010; Chong and Teow, 2013; Shang et al., 2014; Zhuang et al., 2019), which has been termed as Naive-dynamic (ND). Alternatively, more sophisticated techniques have been devised to reduce computational overhead by identifying a smaller subset of the graph affected by changes. These techniques include updating only changed vertices (Aktunc et al., 2015; Yin et al., 2016), processing vertices in the proximity of updated edges (within a specified threshold distance) (Held et al., 2016), disbanding affected communities into lower-level networks (Cordeiro et al., 2016), or employing a dynamic modularity metric to recompute community memberships from scratch (Meng et al., 2016).
Zarayeneh et al. (Zarayeneh and Kalyanaraman, 2021) propose the Delta-screening (DS) approach for updating communities in a dynamic graph. This technique examines edge deletions and insertions to the original graph, and identifies a subset of vertices that are likely to be impacted by the change using the modularity objective. Subsequently, only the identified subsets are processed for community state updates, using the Louvain and Smart Local-Moving (SLM) algorithms (Waltman and Eck, 2013). Recently, Sahu et al. (Sahu et al., 2024b) introduced the parallel Dynamic Frontier (DF) approach for the Louvain algorithm. Given a batch update consisting of edge deletions or edge insertions, the DF approach incrementally identifies an approximate set of affected vertices in the graph with a low run time overhead. They also discuss a parallel algorithm for the Delta-screening (DS) approach.
However, to the best of our knowledge, none of the proposed approaches have been extended to the Leiden algorithm.
@@ 3.1. Community detection
Disjoint community detection involves finding a community membership mapping, $C:V\rightarrow\Gamma$, which maps each vertex $i\in V$ to a community ID $c\in\Gamma$, where $\Gamma$ is the set of community-ids. The vertices within a community $c$ are denoted as $V_{c}$, and the community to which a vertex $i$ belongs is denoted as $C_{i}$. Furthermore, we define the neighbors of vertex $i$ belonging to a community $c$ as $J_{i\rightarrow c}=\{j\ |\ j\in J_{i}\ and\ C_{j}=c\}$, the sum of those edge weights as $K_{i\rightarrow c}=\{w_{ij}\ |\ j\in J_{i\rightarrow c}\}$, the sum of edge weights within $c$ as $\sigma_{c}=\sum_{(i,j)\in E\ and\ C_{i}=C_{j}=c}w_{ij}$, and the total edge weight of $c$ as $\Sigma_{c}=\sum_{(i,j)\in E\ and\ C_{i}=c}w_{ij}$ (Zarayeneh and Kalyanaraman, 2021).
@@ 3.2. Modularity
Modularity is a metric used to assess the quality of communities identified by community detection algorithms, which are heuristic-based (Newman, 2004). It ranges from $[-0.5,1]$ (higher is better), and is calculated as the difference between the fraction of edges within communities and the expected fraction of edges if they were distributed randomly (Brandes et al., 2007). The modularity $Q$ of the obtained communities can be calculated using Equation 1. The delta modularity of moving a vertex $i$ from community $d$ to community $c$, denoted as $\Delta Q_{i:d\rightarrow c}$, is determined using Equation 2.
(1) $Q=\sum_{c\in\Gamma}\left[\frac{\sigma_{c}}{2m}-\left(\frac{\Sigma_{c}}{2m}\right)^{2}\right]$
@@ 3.3. Louvain algorithm
The Louvain method is a greedy, modularity optimization based agglomerative algorithm for detecting communities of high quality in a graph. It works in two phases: in the local-moving phase, each vertex $i$ considers moving to a neighboring community $\{C_{j}\ |\ j\in J_{i}\}$ to maximize modularity increase $\Delta Q_{i:C_{i}\rightarrow C_{j}}$. In the aggregation phase, vertices in the same community are merged into super-vertices. These phases constitute a pass, and repeat until modularity gain stops. This produces a hierarchical structure (dendrogram), with the top level representing the final communities.
@@ 3.5. Dynamic approaches
3.5.1. Naive-dynamic (ND) approach (Aynaud and Guillaume, 2010; Chong and Teow, 2013; Shang et al., 2014; Zhuang et al., 2019)
The Naive dynamic approach is a simple method for detecting communities in dynamic networks. In this approach, vertices are assigned to communities based on the previous snapshot of the graph, and all vertices are processed, regardless of the edge deletions and insertions in the batch update (hence the term naive).
3.5.2. Delta-screening (DS) approach (Zarayeneh and Kalyanaraman, 2021)
@@
This is a dynamic community detection approach that utilizes modularity-based scoring to identify an approximate region of the graph where vertices are likely to alter their community membership. In the DS approach, Zarayeneh et al. first sort the batch update consisting of edge deletions $(i,j)\in\Delta^{t-}$ and insertions $(i,j,w)\in\Delta^{t+}$ by their source vertex ID. For edge deletions within the same community, they mark $i$’s neighbors and $j$’s community as affected. For edge insertions across communities, they select the vertex $j^{*}$ with the highest modularity change among all insertions linked to vertex $i$ and mark $i$’s neighbors and $j^{*}$’s community as affected. Edge deletions between different communities and edge insertions within the same community are unlikely to impact community membership and are therefore ignored. Figure 1(a) shows the vertices (and communities) connected to a single source vertex $i$ that are identified as affected by the DS approach, in response to a batch update with both edge deletions and insertions.
3.5.3. Dynamic Frontier (DF) approach (Sahu et al., 2024b)
@@
This approach first initializes each vertex’s community membership to that obtained in the previous graph snapshot. In instances where edge deletions occur between vertices within the same community or edge insertions between vertices in different communities, the DF approach marks the source vertex $i$ as affected. As batch updates are undirected, both endpoints $i$ and $j$ are effectively marked as affected. Edge deletions spanning different communities or edge insertions within the same community are disregarded (as previously mentioned in Section 3.5.2). Furthermore, when a vertex $i$ alters its community membership during the community detection process, all its neighboring vertices $j\in J_{i}$ are marked as affected, while $i$ is marked as unaffected. To minimize unnecessary computation, an affected vertex $i$ is also marked as unaffected even if its community remains unchanged. Thus, the DF approach follows a graph traversal-like process until the vertices’ community assignments converge. Figure 1(b) shows vertices connected to a single source vertex $i$ identified as affected by the DF approach in response to a batch update.
Table 1. Key terminology used in this report.
Term Definition
Phase Steps in Leiden algorithm: local-moving, refinement, and aggregation.
Pass One full cycle of the three phases. Multiple passes are repeated until convergence.
Community bound Community structure identified during the local-moving phase.
Sub-community Smaller groups formed during the refinement phase within community bounds.
(a) Delta-screening (DS)
(b) Dynamic Frontier (DF)
Figure 1. Illustration of Delta-screening (DS) (Zarayeneh and Kalyanaraman, 2021) and Dynamic Frontier (DF) approaches (Sahu et al., 2024b), in the presence of edge deletions and insertions, represented with dotted lines and doubled lines, respectively. Vertices identified as affected (initial) by each approach are highlighted in brown, and entire communities marked as affected are depicted in light brown.